# Top-of-atmosphere radiation fields over the last millennium reconstructed from proxies

### v1

@@ 2.2 Data assimilation using an ensemble Kalman filter
The Kalman filter is a DA algorithm that optimally combines a prior state, such as a forecast, with available proxy observations, weighted by their uncertainties. In an ensemble Kalman filter (EnKF), the prior and posterior distributions of the state are approximated as normal distributions using samples. Multiple flavors of the EnKF exist. Here, we use a serial ensemble square root filter (Whitaker and Hamill, 2002).
@@
where $\bm{\mathsf{K}}$ is the Kalman gain matrix, and $\bm{\mathsf{H}}$ is the forward operator for estimating the observation from the prior. The individual ensemble members are updated such that the posterior covariance (i.e., their spread), estimates the uncertainty consistently with respect to the full Kalman filter. The Kalman gain matrix encodes covariances between the observation and the state, which allows us to estimate any field that covaries with surface temperature. The full equations for determining $\bm{\mathsf{K}}$ and for updating the ensemble members are presented in the Supplemental Information.
@@
Seasonally resolved proxies are assimilated in the season they represent. Assimilating annually resolved proxies is complicated by the fact that the proxy seasonality can span multiple seasons. Our DA strategy in this case is to update all seasons within this time window when the proxy can be estimated by the ensemble (i.e., once the end of the window is reached), similarly to Meng et al. (2025). For example, a proxy with seasonality MAMJJA is assimilated during the JJA step to update the MAM and JJA values. The update to the last season (JJA) then informs future seasons through the LIM forecast, while updates to past seasons are not propagated forward in time. To perform the EnKF update of multiple seasons simultaneously, we use the time-averaging algorithm from Huntley and Hakim (2010). This algorithm uses the time mean over multiple seasons as $\overline{\mathbf{x}}$ in Eq. (2), then adds the deviations of each season around that prior time mean to the posterior time mean.
We perform DA in EOF space, as opposed to the gridded physical space, which is computationally efficient, since the LIM and the EnKF operate on the same vectors and thus no mapping between spaces is necessary. While this means we cannot mitigate sample and model error using covariance localization (e.g., Anderson, 2012), we instead use a large ensemble and an EOF space that is low-dimensional compared to the ensemble size. Moreover, we perform separate reconstructions with three different LIMs to account for model error. The operator $\bm{\mathsf{H}}$ first maps from the EOF space to the temperature at the proxy location, then maps to the proxy value using the PSM. We map all posteriors from EOF space to physical space as a postprocessing step. We also obtain absolute SIC values by anchoring the anomalies to the merged model–satellite 1961–1990 climatology from Cooper et al. (2025), then clip values outside the interval between zero and one.
@@
We repeat the reconstruction procedure 20 times for each LIM using a Monte Carlo approach. In each iteration, we assimilate a random sample of 80% of the available proxy data. This allows us to sample over uncertainty in the proxy error estimates by randomly removing proxies that may have an outsized impact (Tardif et al., 2019). The remaining 20% of proxies in each iteration are used for independent validation of reconstruction skill in the pre-instrumental period. Each of the 20 iterations uses 400 ensemble members. We reduce the full 8000-member ensemble to 400 members per model prior by subsampling 20 members from each iteration. The analysis below is performed on the resulting 1200-member multi-model ensemble derived from reconstructions from all three model priors.
@@ 2.3 Proxy system models
The EnKF requires a PSM that estimates the proxy value from the current state, corresponding to $\bm{\mathsf{H}}\overline{\mathbf{x}}_{b}$ in Eq. (2). We use linear univariate PSMs that are calibrated over the instrumental period. Such statistical PSMs allow us to estimate the proxy error needed for DA from regression residuals while sacrificing little skill compared to process-based PSMs (Dee et al., 2016; Sanchez et al., 2021).
@@
The linear model takes as input the temperature (SAT or SST, depending on the proxy type) at the nearest gridpoint. For seasonally resolved proxies, we fit one PSM for each season. For annually resolved proxies, we objectively determine the proxy seasonality based on the Bayesian information criterion (as in Tardif et al., 2019), then fit a single PSM that takes the mean over those seasons as input. Seasonal DA allows us to explicitly update seasonal temperatures rather than the annual-mean temperature, whereas annual DA can suffer from seasonality biases (Lücke et al., 2021).
@@
The PSMs are calibrated on GISTEMP v4 (Lenssen et al., 2024; GISTEMP Team, 2025) and ERSSTv5 (Huang et al., 2017). We truncate these temperature datasets to the same EOF basis used for each LIM to include representativeness error in the proxy estimate. We remove GISTEMP data before 1900 since there are spatial discontinuities that may affect the calibration. The full calibration procedure is described in the Supplemental Information.
2.4 Proxy dataset
We assimilate the proxies included in the PAGES2k (PAGES 2k Consortium, 2017) and CoralHydro2k (Walter et al., 2023) databases, and add the Palmyra coral record from Dee et al. (2020). Duplicates are removed based on time resolution, geographical location, and time series correlations. We average all subseasonal proxies to seasonal resolution, and remove proxies that have a time resolution longer than one year since we cannot assimilate them easily. Proxies are excluded if their overlap with the instrumental period is too short (<< 25 years) or their calibration correlation is too low (<< 0.05). The resulting proxy network comprises 647 records, most of which derive from tree rings (64%) and corals (28%; Fig. S2).
The proxy network is relatively sparse before around 1500. In fact, there are less than 10 seasonally resolved proxies before 1650. Importantly, we are still able to reconstruct the seasonal climate since annual proxies are sensitive to certain seasons, and the LIM propagates information to proxy-sparse seasons.
@@ 3 Validation of reconstruction skill
We validate our method using a pseudoproxy experiment (PPE) and the real-proxy reconstruction against instrumental datasets. PPEs help to establish an upper bound on reconstruction fidelity in an idealized scenario (Smerdon, 2012), while the instrumental validation assesses the real reconstruction, albeit over a limited time period and against uncertain verification sources.
3.1 Pseudoproxy experiment
@@
PPEs are particularly useful to assess the feasibility of reconstructing new target variables such as TOA radiation. A climate model simulation serves as “truth,” to be reconstructed based on sparse pseudoproxies that mimic the characteristics of the real proxy network. The challenge is analogous to the real problem, which is to reconstruct the full climate fields given only sparse and noisy proxy observations.
For our PPEs, truth consists of the concatenated last-millennium and historical simulations from the MIROC model. We then reconstruct MIROC fields using LIMs trained on the MPI, CESM, and MRI model priors. This constitutes an imperfect-model experiment, which includes error from a mismatch in LIM dynamics and the EOF basis relative to the truth simulation, analogous to the real reconstruction where the true dynamics are unknown. For each real proxy, the pseudoproxy is drawn from the truth simulation using the real PSM with added noise consistent with the estimated error. The pseudoproxies mimic the real proxies in location, temporal availability, seasonal sensitivity, signal-to-noise ratio, and temperature field sensitivity (SAT or SST).
Results show that the global-mean reconstructions for the fields of interest correlate highly with the truth simulation (Fig. 2a–d). For SAT, decadal variability is tracked well, and the truth is generally within the ensemble spread. The modern warming is somewhat underestimated, possibly because patterns of warming are different in the MIROC model compared to the LIMs, or because the EOF basis, which is derived from the pre-industrial period, cannot represent well the 20th-century trend. The reconstruction is also skillful for the TOA radiation fields. Interannual and decadal variability in the surface temperature induces a radiative response, and the volcanic forcing is partitioned well into RSR and OLR fields, although its magnitude is underestimated. In contrast, while the EEI in the 1900s due to anthropogenic forcing is reconstructed well, the partition into RSR and OLR contributions is flawed. This is because the historical period is not included in the LIM training data, and therefore the LIMs are not aware of the greenhouse gas and aerosol forcing of this period.
@@
Figure 2: Comparison of the pseudoproxy imperfect-model reconstruction with the truth simulation. (a–d) Global-mean time series and their correlation coefficients r over 850–1850. The reconstruction (black) and the truth (yellow) are anomalies with a 20-yr low-pass filter. The truth anomalies are relative to 1961–1990, and the reconstruction has been shifted to match the truth in the 850–1850 mean. DA is the mean of the three single-model reconstructions with the MPI, CESM, and MRI model priors. Shading denotes the 5th–95th percentile range. (e–h) Spatial correlation of annual anomalies over 850–1850. Numbers in the top right corners represent global-mean values.
@@
The PPE also demonstrates spatial reconstruction skill over the last millennium (Fig. 2e–h). Correlations with the truth simulation are positive over most of the globe, although they are smaller over the Southern Ocean, which is relatively sparse in proxies. Like for the global mean, annual skill is highest for SAT and lowest for EEI.
@@ 3.2 Instrumental validation of real-proxy reconstruction
Validation against other datasets allows us to assess the real reconstruction. The SAT, SST, and OHC reconstructions are highly correlated with instrumental products and other reconstructions (Figs. S3 and S5a,b). The SIC reconstruction is skillful at annual and seasonal scales when compared to satellite observations (Figs. S4 and S5c–f). Unassimilated proxies (20% withheld for validation), which are predicted from the reconstruction and compared directly to proxy values, correlate significantly, indicating skill over the whole last millennium (Fig. S6).
@@
The Earth Radiation Budget Experiment (ERBE) satellite observations of TOA radiation overlap with our reconstruction over 1985–1999. We validate against the DEEP-C v5 dataset (Liu and Allan, 2022; Liu et al., 2020), which combines these satellite observations with atmospheric reanalyses and atmosphere-only model simulations. We also validate against the ERBE WFOV Edition 4.1 datasets (ERBE Science Team, 2020a, b), which are direct, recently recalibrated observations but lack some spatial and temporal coverage. The statistical significance of correlations is determined using the random-phase test, a nonparametric test that generates surrogate time series with the same spectral density (Ebisuzaki, 1997).
@@
Correlations of the reconstructed global-mean TOA radiation with the DEEP-C and ERBE data reflect generally skillful tracking of interannual variability (Fig. 3a–c). Much of the skill comes from the tropics and subtropics, but large regions in the higher latitudes have positive correlations as well (Fig. 3d–f). The reconstructed seasonal TOA radiation is skillful as well, but with correlations slightly less than the annual values (Fig. S7). Compared to the PPE, the correlations of the global mean are similar, but the skill mostly comes from the tropics, while it is more uniform in the PPE. The correlations are not always statistically significant and sensitive to small variations because of the short overlap period and high temporal autocorrelation.
@@
Figure 3: Annual TOA radiation anomalies from our reconstruction, the DEEP-C combined product, and ERBE satellite measurements. The anomalies are relative to 1985–1999. Shading denotes the 5th–95th percentile range. (a–c) Global-mean time series and their correlation coefficients r. The p-values indicate if r is different from zero. (d–f) Spatial correlation with DEEP-C. Numbers in the top right corners represent global-mean values.
@@
Our reconstruction also tracks well decadal variations in the upper-ocean heat content due to volcanic eruptions (Fig. 4). The timescale and amplitude of OHC300 changes agree closely with the Cheng et al. (2017) instrumental dataset, which is derived from in-situ temperature profiles. The recovery from the 1963 Agung eruption took twenty years, possibly prolonged by the 1968 Fernandina eruption. We conclude that the reconstruction is generally skillful, despite the spatiotemporal sparseness of the proxy network and the complex relationship between surface temperature and TOA radiation.
@@
Figure 4: Annual upper-ocean heat content from our reconstruction and the Cheng et al. (2017) instrumental dataset. The linear trend over 1960–2000 has been removed to emphasize interannual to decadal variations as in Church et al. (2005). Stratospheric aerosol optical depth (SAOD) at 550 nm is from the CMIP7 historical volcanic forcing dataset (Aubry and CMIP Climate Forcing Task Team, 2025).

### v2

@@ 2.2 Data assimilation using an ensemble Kalman filter
The Kalman filter is a DA algorithm that optimally combines a prior state, such as a forecast, with available proxy observations, weighted by their uncertainties. In an EnKF, the prior and posterior distributions of the state are approximated as normal distributions using samples (Evensen, 1994). Multiple flavors of the EnKF exist. Here, we use a serial ensemble square root filter (Whitaker and Hamill, 2002).
@@
where $\bm{\mathsf{K}}$ is the Kalman gain matrix, and $\bm{\mathsf{H}}$ is the forward operator for estimating the observation from the prior. The individual ensemble members are updated such that the posterior covariance (i.e., their spread), estimates the uncertainty consistently with respect to the non-ensemble Kalman filter. The Kalman gain matrix encodes covariances between the observation and the state, which allows us to estimate any field that covaries with surface temperature. The full equations for determining $\bm{\mathsf{K}}$ and for updating the ensemble members are presented in the Supplemental Information Text S2.
@@
Seasonally resolved proxies are assimilated in the season they represent. Assimilating annually resolved proxies is complicated by the fact that the proxy seasonality can span multiple seasons. Our DA strategy in this case is to update all seasons within this time window when the proxy can be estimated by the ensemble (i.e., once the end of the window is reached), similarly to Meng et al. (2025). For example, a proxy with seasonality MAMJJA is assimilated during the JJA step to update the MAM and JJA values. The update to the last season of the window (JJA) then informs future seasons through the LIM forecast, while updates to past seasons are not propagated forward in time. To perform the EnKF update of multiple seasons simultaneously, we use the time-averaging algorithm from Huntley and Hakim (2010). This algorithm uses the time mean over multiple seasons as $\overline{\mathbf{x}}$ in Eq. (3), then adds the deviations of each season around that prior time mean to the posterior time mean.
We perform DA directly on the state vector (i.e., in EOF space) rather than in physical space. This improves computational and storage efficiency because both LIM and EnKF operate in the same low-dimensional EOF space. However, operating in EOF space prevents the use of covariance localization (e.g., Anderson, 2012). Instead, we mitigate sample error by using an ensemble that is much larger than the low-dimensional state vector, and we account for model error by running separate reconstructions with five distinct LIMs. To estimate the proxy value in Eq. (3), the observation operator $\bm{\mathsf{H}}$ must map the state vector to the physical temperature at the proxy location, and subsequently to the proxy value using the PSM (e.g., Hakim et al., 2022).
SIC is physically bounded between zero and one, but the Gaussian distributions of the EnKF do not enforce this constraint. To obtain valid values, we convert SIC anomalies to absolute values by anchoring them to a 1961–1990 climatology, then clip SIC values falling outside the $[0,1]$ interval. The climatology, obtained from Cooper et al. (2025), is based on the multi-model mean from eight CMIP6 historical simulations before 1979, and on satellite observations thereafter.
@@
The LIMs from each of the five model priors are used separately in the DA algorithm, allowing us to sample structural error in the training data, such as differences in covariance relationships. We repeat the reconstruction procedure 20 times for each LIM using a Monte Carlo approach. In each iteration, we assimilate a random sample of 80% of the available proxy data. This allows us to sample over uncertainty in the proxy error estimates by randomly removing proxies that may have an outsized impact (Tardif et al., 2019). The remaining 20% of proxies in each iteration are used for independent validation of reconstruction skill in the pre-instrumental period. Each of the 20 iterations uses 400 ensemble members. We reduce the full 8000-member ensemble per model prior to 400 members by subsampling 20 members from each iteration. The analysis below is performed on the resulting 2000-member multi-model ensemble derived from reconstructions with all five model priors.
@@ 2.3 Proxy system models
The EnKF requires PSMs that estimate the proxy values from the current state, corresponding to $\bm{\mathsf{H}}\overline{\mathbf{x}}_{b}$ in Eq. (3). We use linear univariate PSMs that are calibrated over the instrumental period. Such statistical PSMs allow for estimating the proxy error needed for DA from regression residuals, while sacrificing little skill compared to process-based PSMs for most proxy types (Dee et al., 2016; Sanchez et al., 2025). Tree-ring width (TRW) proxies, however, may be poorly modeled by a univariate regression in temperature due to the confounding influence of moisture (Dee et al., 2016). We address this by removing TRW proxies that are primarily moisture-sensitive.
@@
The linear model takes as input the temperature (SAT for terrestrial, SST for marine proxies) at the nearest gridpoint. For seasonally resolved proxies, we fit one PSM for each season. For annually resolved proxies, we objectively determine the proxy seasonality based on the Bayesian information criterion (as in Tardif et al., 2019), then fit a single PSM that takes the mean over those seasons as input. Seasonal DA allows us to explicitly update seasonal temperatures rather than the annual-mean temperature, whereas annual DA can suffer from seasonality biases (Lücke et al., 2021).
@@
The PSMs are calibrated on GISTEMP v4 (Lenssen et al., 2024; GISTEMP Team, 2025) and ERSSTv6 (Huang et al., 2025a, b). We truncate these temperature datasets to the same EOF basis used for each LIM to include representativeness error in the estimated observation error. We remove GISTEMP data before 1900 since there are spatial discontinuities that may affect the calibration. We also remove calibration data after 2000 to avoid the divergence problem, i.e., the decoupling of proxy values from local temperature, such as is common for tree-ring proxies (e.g., D’Arrigo et al., 2008). The full calibration procedure is described in the Supplemental Information Text S3.
2.4 Proxy network
[…]
We average all subseasonal proxies to seasonal resolution, and remove proxies that have a time resolution longer than one year since we cannot assimilate them easily. Proxies are excluded if their overlap with the calibration period is too short (<< 25 years), their calibration correlation is too low (<0.10<0.10), or their error autocorrelation is too high (>0.90>0.90; errors are assumed uncorrelated in time in the Kalman filter). The resulting proxy network comprises 382 records, most of which derive from tree rings (58%) and corals (30%). The proxies have broad spatial coverage, but few are located in the ocean interior, particularly outside the tropics (Fig. S5). The proxy network becomes increasingly sparse toward the beginning of the last millennium, and there are only four seasonally resolved proxies before 1500. Importantly, since annual proxies are sensitive to specific seasons, and the LIM propagates information through proxy-sparse seasons, the algorithm yields seasonal climate information even when there are no seasonal proxies.
2.5 Skill metrics
We validate the reconstruction skill in pseudoproxy experiments, against instrumental datasets, and against withheld proxies. Our primary validation metrics are the Pearson correlation coefficient $r$ and the coefficient of efficiency (Nash and Sutcliffe, 1970)
$\text{CE}=1-\frac{\sum_{i}(v_{i}-x_{i})^{2}}{\sum_{i}(v_{i}-\bar{v})^{2}},$
where $x_{i}$ is the reconstruction time series, $v_{i}$ is the verification time series, and $\bar{v}$ is the time mean of $v_{i}$. Compared to $r$, the CE thus also accounts for differences in signal amplitude and mean, measured relative to the verification variability.
The statistical significance of correlations is determined using the random-phase test, a nonparametric test that generates surrogate time series with the same spectral density (Ebisuzaki, 1997). This accounts for spectral properties such as autocorrelation, which reduces the effective sample size (e.g., Bretherton et al., 1999). For p-values, we use a significance level of $\alpha=0.05$ with the null hypothesis that $r=0$. The alternative hypothesis is $r>0$ if reconstruction skill is compared, and $r\neq 0$ for physical correlations between variables. In some figures and tables, we denote highly significant correlations ($p\leq 0.01$) with **, significant correlations ($p\leq 0.05$) with *, and non-significant correlations ($p>0.05$) with n.s.
@@ 3 Validation of reconstruction skill
We validate our method using pseudoproxy experiments (PPEs) and compare the real-proxy reconstruction against instrumental datasets. PPEs help to establish an upper bound on reconstruction fidelity in an idealized scenario (Smerdon, 2012), while instrumental validation assesses the real reconstruction, albeit over a limited time period and against uncertain verification sources (e.g., Thorne et al., 2026; Chan et al., 2025).
3.1 Realistic pseudoproxy experiment
@@
PPEs are particularly useful to assess the feasibility of reconstructing new target variables such as TOA radiation. A climate model simulation serves as “truth,” sampled only by sparse and noisy pseudoproxies that mimic the characteristics of the real proxy network. The challenge is analogous to the real problem, with the benefit that skill can be easily quantified (Smerdon, 2012; Steiger et al., 2014).
@@
Figure 2: Comparison of the pseudoproxy imperfect-model reconstruction with the truth simulation. (a–d) Global-mean time series and their correlation coefficients r over 850–1850. The reconstruction (black) and the truth (yellow) are anomalies relative to 850–1850 with a 20-yr low-pass filter. Radiation after 1850 is dashed since it includes anthropogenic forcing, which we cannot reconstruct. DA is the mean of the four single-model reconstructions with the MPI, CESM, MRI, and EC-Earth model priors. Carets indicate volcanic eruptions, scaled by their volcanic stratospheric sulfur injection, from the eVolv2k (Toohey and Sigl, 2017) and the CMIP7 historical volcanic forcing datasets (Aubry et al., 2025; small eruptions with a VSSI below $0.5\,\mathrm{Tg\ S}$ removed). All correlations of global means are significant as denoted by (**). Shading denotes the 5th–95th percentile range. (e–h) Spatial correlation of annual anomalies over 850–1850 between reconstruction and truth. Correlations that are not significantly positive ($\alpha=0.05$) are hatched. Numbers in the top right corners are global means of the spatial correlation.
For our realistic PPEs, truth consists of the concatenated last-millennium and historical simulations from the MIROC model. We then reconstruct MIROC fields using LIMs trained on the MPI, CESM, MRI, and EC-Earth model priors. This constitutes an imperfect-model experiment, which includes error from a mismatch in LIM dynamics and the EOF basis relative to the truth simulation, analogous to the real reconstruction where the true dynamics are unknown. For each real proxy, a pseudoproxy is drawn from the truth simulation using the real PSM with added noise consistent with the estimated error. The pseudoproxies mimic the real proxies in location, temporal availability, seasonal sensitivity, signal-to-noise ratio, and temperature field sensitivity (SAT or SST).
Results show that the global-mean reconstructions for the fields of interest correlate significantly with the truth simulation (Fig. 2a–d). For SAT, decadal variability is tracked well (annual $r=0.8$), and the truth is generally within the ensemble spread. Modern warming is somewhat underestimated, possibly because patterns of warming are different in the MIROC model compared to the LIMs (Fig. 2a). The reconstruction is also mostly skillful for the TOA radiation fields, although correlations and CE are lower for EEI than for RSR and OLR. Large excursions in TOA radiation coincide with major volcanic eruptions, although the magnitude is underestimated, particularly in the EEI. The variance of the EEI in the PPE is similar to the real reconstruction, suggesting that the PPE may be representative of the real problem. Arctic SIC validates well on sea-ice area and spatial patterns, in contrast to Antarctic SIC, which shows low correlations and a large spread across model priors (Fig. S3). Reconstruction skill for all fields is significantly reduced in the PPE before 1100, when fewer than 70 proxies are available, none of them in the tropical Pacific.
@@
The PPE also demonstrates spatial reconstruction skill over the last millennium (Fig. 2e–h). Correlations with the truth simulation are positive in most regions, although they are smaller over the Southern Ocean, which is relatively sparse in proxies. Like for the global mean, annual skill is highest for SAT and lowest for EEI.
3.2 Pseudoproxy experiment for forcing and response
We run further PPEs to test if our DA system is able to reconstruct both the response and the natural forcing, despite their distinct lead/lag relationships with temperature (e.g., Proistosescu et al., 2018). In these PPEs, we separate the forcing and response in the state vector. The truth and training data come from CMIP6 “hist-nat” simulations (Gillett et al., 2016), with forcing derived from “piClim-histnat” simulations (Pincus et al., 2016), of the CanESM5 (Swart et al., 2019) and NorESM2-LM (Seland et al., 2020) models.
We find skillful reconstructions of the response, natural forcing, and 300-m OHC (Fig. S4, Table S1). Correlations are particularly high for the response and OHC. While forcing skill is relatively high in perfect-model experiments, it is not always significant in imperfect-model experiments. The magnitude of volcanic forcing is greatly underestimated, by up to a factor of five, similar to the realistic PPE experiment (Fig. 2b). The lead/lag relationships of the response and forcing in the models are reproduced by the reconstruction (Fig. S4d), which is critical for proxy information to inform the energy imbalance through DA. Natural forcing also includes solar variability, although its magnitude is below the noise level of the reconstruction. We conclude that reconstructing both response and natural forcing is possible, but with a much reduced magnitude for volcanic forcing. We hypothesize that OHC plays a critical role in linking the energy budget and radiation to temperatures through proxies.
@@ 3.3 Instrumental validation of real-proxy reconstruction
Validation against independent datasets allows us to assess the real reconstruction. The SAT, SST, and OHC reconstructions are highly correlated with instrumental products and other reconstructions (Figs. S7 and S9a,b). The Arctic SIC reconstruction is skillful at annual and seasonal scales when compared to satellite observations (Figs. S8 and S9c–f), but Antarctic SIC is not. Further, unassimilated proxies (20% withheld for validation), which are predicted from the reconstruction and compared directly to proxy values, correlate significantly in the network mean, indicating skill over the whole last millennium (Fig. S6).
@@
The Earth Radiation Budget Experiment (ERBE) satellite observations of TOA radiation overlap with our reconstruction over 1985–1999. We validate against the DEEP-C v5 dataset (Liu et al., 2020; Liu and Allan, 2022), which combines these satellite observations with atmospheric reanalyses and atmosphere-only model simulations. We also validate against the ERBE WFOV Edition 4.1 datasets (ERBE Science Team, 2020), which are direct, recently recalibrated observations but lack complete spatial and temporal coverage.
@@
Reconstructed global-mean TOA radiation qualitatively tracks interannual variability in the DEEP-C and ERBE data, with positive but not always significant correlations of $r>0.4$ (Fig. 3a–c). Most of this variability represents the radiative response to internal variability, but there is a forced signal from the 1991 Mt. Pinatubo eruption. As in the PPEs (Section 33.1,3.2), this volcanic forcing is underestimated. Much of the skill comes from the tropics and subtropics, likely mediated by ENSO, but large regions in the mid-latitudes and polar regions also have positive correlations (Fig. 3d–f). The reduced skill at higher latitudes is likely the result of a weaker temperature–radiation relationship, mainly due to weather noise from extratropical storms in TOA radiation (e.g., Trenberth et al., 2015). The reconstructed seasonal TOA radiation is also skillful, but with correlations slightly less than the annual values (Fig. S10). Compared to the PPE, the correlations of the global mean are similar, but the skill mostly comes from the tropics, while it is more uniform in the PPE. The correlations are not robust due to the short overlap period and high temporal autocorrelation.
@@
Figure 3: Annual TOA radiation anomalies from our reconstruction, the DEEP-C combined product, and ERBE satellite measurements. The linear trend over 1985–1999 has been removed. Shading denotes the 5th–95th percentile range. (a–c) Global-mean time series and their correlation coefficients r. (d–f) Spatial correlation of reconstruction with DEEP-C. Numbers in the top right corners are global-mean values. Significance of spatial correlations is not indicated since most are non-significant due to the small effective sample size ($n_{\text{eff}}\approx 10$ to $15$ for global means).
@@
Our reconstruction also correlates well ($r=0.72$, $p\leq 0.01$) with decadal variations in instrumental estimates of the upper-ocean heat content (Fig. 4). For comparison, we use the Minière et al. (2026) compilation of instrumental datasets, derived primarily using in-situ temperature profiles. Most of the decadal variability in OHC appears forced by volcanic eruptions. The recovery of OHC following the 1963 Agung eruption took multiple decades, possibly prolonged by the 1968 Fernandina eruption, unforced decadal variability, and tropospheric aerosol forcing. We conclude that our reconstruction of the energy budget has many skillful aspects, despite the spatiotemporal sparseness of the proxy network and the complex relationship between surface temperature and TOA radiation.
@@
Figure 4: Annual upper-ocean heat content from our reconstruction and the Minière et al. (2026) compilation of instrumental datasets. The linear trend over 1960–2000 has been removed to emphasize interannual to decadal variations as in Church et al. (2005). Only instrumental datasets that span the full period are used, leaving a compilation of eight. Stratospheric aerosol optical depth (SAOD) at 550 nm is from the CMIP7 historical volcanic forcing dataset (Aubry et al., 2025). Shading denotes the 5th–95th percentile range.